OpenStudy (anonymous):

establish the relation- v^2-u^2=2as, using velocity time graph....

OpenStudy (anonymous):

solve plzz....

OpenStudy (anonymous):

...............................?????????????????

OpenStudy (jamesj):

Ok, so on a graph of velocity vs. time, what does the graph look like with constant acceleration a?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

What kinds of graphs can you have if acceleration a is a constant?

OpenStudy (anonymous):

velocity- time graph

OpenStudy (jamesj):

Yes, the y axis is velocity, and the x-axis is time. What kinds of curves do we have for v as a function of t, v = v(t), if a is a constant? Do we have just constant functions, i.e., horizontal lines? Do we have straight lines in general? Do we have parabolas? Do we have other sorts of graphs?

OpenStudy (anonymous):

plzzzz do in ur way just/....

OpenStudy (anonymous):

????????????????????????????????????????????????

OpenStudy (anonymous):

plzzzzz fst

OpenStudy (anonymous):

@@JamesJ

OpenStudy (jamesj):

Look, can you answer my question?

OpenStudy (jamesj):

If a is a constant and u is the initial velocity, what is the formula for v, the final velocity? v = .... what, as a function of u and t?

OpenStudy (anonymous):

v= u+at

OpenStudy (jamesj):

correct. So in a velocity-time diagram, what kind of curve is that?

OpenStudy (anonymous):

straight curve

OpenStudy (anonymous):

i think so....

OpenStudy (jamesj):

Yes, a straight line. Now, on that diagram, what is s?

OpenStudy (anonymous):

distance

OpenStudy (jamesj):

The area under the velocity curve is the displacement, s, as indicated in the diagram I just posted. Now, given that diagram, you will want to write down a formula for the area, s, as a function of u, v and a.

OpenStudy (jamesj):

You already know that you want v^2 - u^2 = 2as, hence it must be that \[ s =\frac{v^2 - u^2}{2a}\] You will need to convince yourself that that is correct.

OpenStudy (jamesj):

Hint, because a = (v-u)/t, we also know that t = (v-u)/a.

OpenStudy (anonymous):

u sure that it was right..??????

OpenStudy (anonymous):

????????????????

OpenStudy (jamesj):

Absolutely. What is the formula for the area under that curve, just working from first principles?

OpenStudy (jamesj):

Consider it as a rectangle and a triangle.

OpenStudy (anonymous):

half*b*h

OpenStudy (jamesj):

I'm going to post one more diagram for you ... one sec

OpenStudy (anonymous):

okk

OpenStudy (jamesj):

The area under the curve has two parts: a rectangle and a triangle. We build up the area of the total by finding the area of each piece. That is clearly indicated on this diagram.

OpenStudy (anonymous):

confused....

OpenStudy (jamesj):

What are you confused about?